# Starfish lemma via birational quasi-isomorphisms

### v2

@@ 1 Introduction
Cluster algebras were introduced by S. Fomin and A. Zelevinsky in [8]. Since then, the subject has developed in numerous directions with applications in mathematical physics, Poisson geometry, combinatorics, representation theory, etc. We refer the reader to [6, 12, 22].
Let us briefly recall the definition of a cluster structure of geometric type. Let $R$ be an integral domain. One starts with a collection $\mathbf{x}:=(x_{1},x_{2},\ldots,x_{N+M})$ of algebraically independent elements of $\operatorname{Frac}R$ over $\mathbb{Q}$ and an integer-valued matrix $B:=(b_{ij}\ |\ i\in[1,N],\,j\in[1,N+M])$ such that $B^{[1,N]}$ (the first $N$ columns, the principal part) is skew-symmetric. The collection $\mathbf{x}$ is called an extended cluster, its elements $x_{1},x_{2},\ldots,x_{N}$ are called cluster variables, $x_{N+1},\ldots,x_{N+M}$ are called frozen variables; in addition, $B$ is called an exchange matrix and the pair $(\mathbf{x},B)$ is called an extended seed. One can replace a cluster variable $x_{k}$ in $\mathbf{x}$ with another element $x_{k}^{\prime}\in\operatorname{Frac}R$ via the mutation relation:
As a result, one obtains a new extended cluster $\mathbf{x}^{\prime}_{k}:=(\mathbf{x}\setminus\{x_{k}\})\cup\{x_{k}^{\prime}\}$. The matrix $B$ also changes with each mutation according to some pattern. Fixing an extended seed $\Sigma_{0}:=(\mathbf{x}_{0},B_{0})$ (the initial extended seed), one considers all possible extended seeds that can be obtained from $\Sigma_{0}$ via iterative applications of mutations. The resulting collection $\mathcal{C}:=\mathcal{C}(\Sigma_{0})$ of extended seeds is called a cluster structure of geometric type in $R$. A cluster algebra $\mathcal{A}(\mathcal{C})$ is the $\mathbb{Z}$-subalgebra of $\operatorname{Frac}R$ generated by all cluster and frozen variables in $\mathcal{C}$. The rank of a cluster algebra is the number $N$ of cluster variables in a single (extended) cluster. An upper cluster algebra $\bar{\mathcal{A}}(\mathcal{C})$ is the intersection of all Laurent subalgebras $\mathcal{L}(\mathbf{x}):=\mathbb{Z}[x_{1}^{\pm 1},x_{2}^{\pm 1},\ldots,x_{N}^{\pm 1},x_{N+1},\ldots,x_{N+M}],\ \mathbf{x}\in\mathcal{C}.$ A central result in cluster theory is the Laurent phenomenon, which states that there is an inclusion $\mathcal{A}(\mathcal{C})\subseteq\bar{\mathcal{A}}(\mathcal{C})$. If $R$ is an algebra over a field $\mathbb{K}$, one typically extends scalars for the above algebras: $\mathcal{A}_{\mathbb{K}}(\mathcal{C}):=\mathcal{A}(\mathcal{C})\otimes\mathbb{K}$, $\bar{\mathcal{A}}_{\mathbb{K}}(\mathcal{C})=\bar{\mathcal{A}}(\mathcal{C})\otimes\mathbb{K}$, $\mathcal{L}_{\mathbb{K}}(\mathbf{x}):=\mathcal{L}(\mathbf{x})\otimes\mathbb{K}$, etc.
One of the objectives of S. Fomin and A. Zelevinsky’s program was identifying coordinate rings $\mathbb{K}[V]$ of interesting varieties $V$ over a field $\mathbb{K}$ in Lie theory as cluster algebras; in other words, finding a cluster structure $\mathcal{C}$ in $\mathbb{K}[V]$ such that $\mathcal{A}_{\mathbb{K}}(\mathcal{C})=\mathbb{K}[V]$. It was realized in [2] that the equality $\bar{\mathcal{A}}_{\mathbb{K}}(\mathcal{C})=\mathbb{K}[V]$ is also of interest in cluster theory. We call a cluster structure $\mathcal{C}$ in $\mathbb{K}[V]$ complete if $\bar{\mathcal{A}}_{\mathbb{K}}(\mathcal{C})=\mathbb{K}[V]$. In order to deduce completeness, one typically invokes the following two results: the Starfish lemma [2, Theorem 2.10] and a result on upper bounds [2, Corollary 1.9].
Let us call a variable $x$ regular if $x\in R$, and two variables $x,y\in R$ coprime if they are not contained in the same prime ideal of height $1$ of $R$. We call an integral domain $R$ normal if $R$ is integrally closed in its field of fractions.
Proposition 1.1.
@@
    All variables in $\mathbf{x}_{0}$ are regular;
@@
    All cluster variables in $\mathbf{x}_{0}$ are pairwise coprime;
@@
    For each cluster variable $x$ in $\mathbf{x}_{0}$, $x^{\prime}$ is regular and coprime with $x$ in $R$.
Then (Starfish lemma) $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$. If in addition,
  1. 3)
    The initial exchange matrix $B_{0}$ has full rank;
  2. 4)
    There are inclusions $\mathcal{L}(\mathbf{x}_{0})\supseteq R$ and $\mathcal{L}(\mathbf{x}_{i}^{\prime})\supseteq R$ for each mutation $\mathbf{x}_{i}^{\prime}$ of $\mathbf{x}_{0}$, $i\in[1,N]$.
Then (Result on Upper bounds and Starfish lemma) $\bar{\mathcal{A}}(\mathcal{C})=R$.
The goal of this paper is as follows. Let $\tilde{\mathcal{C}}$ be a cluster structure in a normal Noetherian domain $\tilde{R}$ that satisfies the conditions of Proposition 1.1. Introduce all the data for $\tilde{\mathcal{C}}$ as for $\mathcal{C}$, decorating it with a tilde (for instance, $\tilde{N}$ is the rank of $\tilde{\mathcal{C}}$). Assume that $\tilde{N}+\tilde{M}=N+M$ and $\tilde{N}<N$ (for instance, $\mathcal{C}$ could be obtained by unfreezing some frozen variables of $\tilde{\mathcal{C}}$). Our goal is to design tools for verifying the conditions of Proposition 1.1 for $\mathcal{C}$ by comparing it with $\tilde{\mathcal{C}}$ via a quasi-isomorphism, and if $\tilde{\mathcal{C}}$ is complete, we would like to claim that $\mathcal{C}$ is complete as well.
One notion of a quasi-isomorphism was introduced by C. Fraser in [9]. In Fraser’s setting, $N=\tilde{N}$, and $M$ and $\tilde{M}$ can be arbitarry. A quasi-isomorphism is a field isomorphism equivariant with respect to mutations; it sets up a correspondence between cluster variables in $\mathbf{x}_{0}$ and $\tilde{\mathbf{x}}_{0}$, and it sends a cluster variable from $\mathbf{x}_{0}$ to a cluster variable in $\tilde{\mathbf{x}}_{0}$ times a monomial in frozen variables.
In our setting, $\tilde{N}<N$ and $\tilde{N}+\tilde{M}=N+M$, and some cluster variables in $\mathbf{x}_{0}$ correspond to frozen variables in $\tilde{\mathbf{x}}_{0}$; we call such variables marked, and the quasi-isomorphism is required to send a variable from $\mathbf{x}_{0}$ (whether it’s cluster or frozen) to the corresponding variable in $\tilde{\mathbf{x}}_{0}$ times a monomial in the marked variables. If one freezes the marked variables in ${\mathbf{x}}_{0}$, one recovers C. Fraser’s notion. However, one key difference is that we require the quasi-isomorphism to be an isomorphism of the rings $R$ and $\tilde{R}$ localized at the marked variables. In this way, we are able to transfer algebraic properties (such as regularity, coprimality, etc.) of cluster and frozen variables from $\tilde{R}$ to $R$.
Our original motivation comes from the Gekhtman-Shapiro-Vainshtein (GSV) program on constructing generalized cluster structures in the coordinate rings $\mathbb{C}[G]$ of connected simply connected simple complex algebraic groups $G$ [12, 14]. The coordinate ring $\mathbb{C}[G]$ can support several cluster structures of different ranks, and the techniques presented in this paper allow one to deduce the completeness of a cluster structure of a larger rank from the properties of cluster structures of smaller ranks. Examples of birational quasi-isomorphisms subject to our requirements appeared in in [11, 17, 21], and early examples appeared in [10].
One can view our work as a step towards solving the following fundamental problem: Given a complete cluster structure $\mathcal{C}$ in a normal Noetherian domain $R$, how to unfreeze a given variable so that the new cluster structure is still complete? At the abstract level of cluster algebra, unfreezing is a trivial operation; however, it is nontrivial if one insists on preserving the relation $\bar{\mathcal{A}}(\mathcal{C})=R$. We provide some examples in Appendix A, where one can see that unfreezing requires changing the realization of cluster and frozen variables as elements of the ring $R$.
The paper is organized as follows. In Section 2, we provide an expanded account on the theory of cluster algebras of geometric type. In Section 3, we provide a background on quasi-isomorphisms subject to our conditions. In Section 4, we study the case of a single quasi-isomorphism (that is, we compare $\mathcal{C}$ only with one $\tilde{\mathcal{C}}$); the main result in Section 4 is Proposition 4.9. In Section 5, we study the case of complementary birational quasi-isomorphisms (that is, we compare $\mathcal{C}$ with a pair of cluster structures $\tilde{\mathcal{C}}$ and $\hat{\mathcal{C}}$); therein, the main result is Proposition 5.4. In Appendix A, we provide explicit examples of cluster structures in $\mathbb{C}[\operatorname{SL}_{3}]$ along with birational quasi-isomorphisms.
Acknowledgements.
The author would like to thank M. Gekhtman for reading the early draft of the paper and for comments and suggestions. This work was supported by the Institute for Basic Science (IBS-R003-D1).
@@ 2 Background on cluster algebras
In this section, we provide an expanded account on the theory of cluster algebras of geometric type. The main references are [6, 7, 12] and the original papers [2, 8]. For this section, let us fix numbers $N$ and $M$ and an ambient field $\mathcal{F}$, which is the field of rational functions in $N+M$ variables with coefficients in $\mathbb{Q}$.
@@ 3.2 Quasi-isomorphisms
    Given the $y$-variables $y_{1},\ldots,y_{N}$ that correspond to $\mathbf{x}$, $\mathcal{Q}(y_{i})=\tilde{y}_{\kappa(i)}$;
@@
$\Lambda_{i}(\mathbf{x}):=\begin{cases}\dfrac{w_{i}^{\prime}-\tilde{w}_{\kappa(i)}^{\prime}}{\tilde{w}_{\kappa(m)}}\ &\text{if}\ i\neq m\\[3.0pt] 1\ &\text{if}\ i=m.\end{cases}$ (3.12)
@@
4 Single birational quasi-isomorphism
Let $R$ and $\tilde{R}$ be normal Noetherian domains. Assume that $(\mathcal{C},R)$ and $(\tilde{\mathcal{C}},\tilde{R})$ are cluster structures of geometric type in $R$ and $\tilde{R}$ with fixed initial extended seeds $(\mathbf{x}_{0},B_{0})$ and $(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$, and such that $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ is a related triple. We also assume there is a quasi-isomorphism $\mathcal{Q}:\mathcal{C}\rightarrow\tilde{\mathcal{C}}$. The objective of this section is to deduce the conditions of Proposition 1.1 for $(\mathcal{C},R)$ assuming the conditions hold for $(\tilde{\mathcal{C}},\tilde{R})$. The main result of this section is Proposition 4.9.
@@ 4 Single birational quasi-isomorphism
The quasi-isomorphism $\mathcal{Q}:\mathcal{C}\rightarrow\tilde{\mathcal{C}}$ is called birational if the isomorphism [^6] By $\mathbf{x}_{0}(\kappa)^{\pm 1}$ we mean the localization at each marked variable $x_{i}\in\mathbf{x}_{0}(\kappa)$.
@@
For the rest of the section, we assume that $\mathcal{Q}$ is a birational quasi-isomorphism.

### v3

@@ 1 Introduction
Cluster algebras were introduced by S. Fomin and A. Zelevinsky in [8]. The main objective of the theory is to represent the coordinate ring of a complex variety $V$ (often arising in Lie theory) in the form $\mathcal{A}_{\mathbb{C}}(\mathcal{C})=\mathbb{C}[V]$ where $\mathcal{C}$ is a cluster structure in $\mathbb{C}[V]$ and $\mathcal{A}_{\mathbb{C}}(\mathcal{C})$ is a cluster algebra. More generally, one aims to prove the equality $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})=\mathbb{C}[V]$ where $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})$ is the upper cluster algebra.
Many examples of cluster algebras are known. Early examples include the Grassmannian [22], double Bruhat cells [2] and simply connected simple complex algebraic groups [14]. More recent examples include open Richardson varieties [19] and Bott-Samelson cells [23].
The main technique for proving the inclusion $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})\subseteq\mathbb{C}[V]$ (or $\mathcal{A}_{\mathbb{C}}(\mathcal{C})\subseteq\mathbb{C}[V]$) is the so-called Starfish lemma (see Proposition 2.9). The lemma requires verifying certain regularity and coprimality conditions in the initial cluster and its one-step mutations. To show the reverse inclusion $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})\supseteq\mathbb{C}[V]$, one typically applies a result on upper bounds (see Proposition 2.7).
The work of M. Gekhtman, M. Shapiro and A. Vainshtein shows that the coordinate ring of a simply connected simple complex algebraic group $G$ can support several cluster structures of different ranks [10, 14, 16]. In general, it is harder to verify the conditions of the Starfish lemma in $\mathbb{C}[G]$ for cluster structures of higher ranks. However, as was shown in [10, 11, 17, 24], such cluster structures are typically related by a birational quasi-isomorphism.
In this paper, we axiomatize the notion of a birational quasi-isomorphism, and use it to prove an analogue of the Starfish lemma. Specifically, given cluster structures $\tilde{\mathcal{C}}$ and $\mathcal{C}$ of ranks $\tilde{N}<N$ in normal Noetherian domains $\tilde{R}$ and $R$, we assume the conditions of the Starfish lemma hold for $\tilde{\mathcal{C}}$; then, if there is a birational quasi-isomorphism between $(\mathcal{C},R)$ and $(\tilde{\mathcal{C}},\tilde{R})$, we deduce the corresponding conditions for $(\mathcal{C},R)$. In this setting, we only need to verify conditions for the so-called marked variables (see Definition 3.2; these variables are frozen in $\tilde{\mathcal{C}}$ and cluster in $\mathcal{C}$).
Our results are particularly useful in case there is a third cluster structure $(\hat{\mathcal{C}},\hat{R})$ with $\hat{N}<N$ that is birationally quasi-isomorphic to $(\mathcal{C},R)$. In this case, the conditions on the marked variables are mild; moreover, if the given rings $R$, $\tilde{R}$ and $\hat{R}$ are UFD’s, the conditions on the marked variables can be dropped. The results of the paper are applied in our recent work [17] (see Appendix A for an explicit example).
The notion of a quasi-isomorphism was previously studied by C. Fraser in [9]. The notion that we need for the Starfish lemma (and for covering the existing examples) is different from the Fraser’s notion in the following ways:
@@
    Fraser’s definition requires cluster structures $\mathcal{C}$ and $\tilde{\mathcal{C}}$ to have equal rank $\tilde{N}=N$, while allowing arbitrary coefficients (in particular, any numbers of frozen variables). In our case, we require $\tilde{N}<N$, but the total number of variables (cluster plus frozen) in a single extended cluster has to be the same.
@@
    In Fraser’s setup, a quasi-isomorphism sends a variable from $\tilde{\mathcal{C}}$ to a variable from $\mathcal{C}$ times a monomial in frozen variables. In our case, the monomial has to be in the marked variables.
@@
    Finally, we require the quasi-isomorphism to be an isomorphism between the localizations of $R$ and $\tilde{R}$ at the marked variables.
The paper is organized as follows. In Section 2, we review the necessary background on cluster algebras. In Section 3, we review the notion of a quasi-isomorphism adjusted for our purposes. In Section 4, we state our main results. In Section 5 and Section 6, we provide proofs of the main results along with results of lesser importance. In Appendix A, we illustrate our theory in $\mathbb{C}[\operatorname{SL}_{3}]$.
Acknowledgments.
The author would like to thank M. Gekhtman for reading the early draft of the paper, for comments and suggestions. This work was supported by the Institute for Basic Science (IBS-R003-D1).
@@ 2 Background on cluster algebras
In this section, we provide a background on the theory of cluster algebras of geometric type. The main references are [6, 7, 12] and the original papers [2, 8]. For this section, let us fix numbers $N$ and $M$ and an ambient field $\mathcal{F}$, which is the field of rational functions in $N+M$ variables with coefficients in $\mathbb{Q}$.
@@ 3.2 Quasi-isomorphisms
    For $i\in[1,N]\setminus\mathcal{I}(\kappa)$ and the $y$-variables corresponding to $\mathbf{x}$, $\mathcal{Q}(y_{i})=\tilde{y}_{\kappa(i)}$;
@@
$\Lambda_{i}(\mathbf{x})=\begin{cases}\dfrac{w_{i}^{\prime}-\tilde{w}_{\kappa(i)}^{\prime}}{\tilde{w}_{\kappa(m)}}\ &\text{if}\ i\neq m\\[3.0pt] 1\ &\text{if}\ i=m.\end{cases}$ (3.12)
@@
3.3 Birational quasi-isomorphisms
We continue in the setup of the previous section. We assume further that $\mathcal{C}$ and $\tilde{\mathcal{C}}$ are cluster structures in normal Noetherian domains $R$ and $\tilde{R}$, respectively.
@@ 3.3 Birational quasi-isomorphisms
A quasi-isomorphism $\mathcal{Q}:\mathcal{C}\rightarrow\tilde{\mathcal{C}}$ is called birational if the marked variables $\mathbf{x}_{0}(\kappa)$ and $\tilde{\mathbf{x}}_{0}(\kappa)$ are regular, and the isomorphism [^6] By $\mathbf{x}_{0}(\kappa)^{\pm 1}$ we mean the localization at each marked variable $x_{i}\in\mathbf{x}_{0}(\kappa)$.
@@
Assume there is another cluster structure $\hat{\mathcal{C}}$ in a normal Noetherian domain $\hat{R}$, and such that $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$ is a related triple. Similarly to the case of $\tilde{\mathcal{C}}$, all the data associated with $\hat{\mathcal{C}}$ is decorated with a hat.
Definition 3.12.
Assume there is a pair $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$ and $\mathcal{Q}^{\prime}:(\mathcal{C},R)\rightarrow(\hat{\mathcal{C}},\hat{R})$ of birational quasi-isomorphisms. Then the pair is called complementary if $\mathcal{I}(\kappa)\cap\mathcal{I}(\kappa^{\prime})=\emptyset$.
4 Main results
In this section, we state the main results of the paper. We begin with a technical definition.
Definition 4.1.
  1. 1)
    All variables in $A$ are regular;
  2. 2)
    All cluster variables in $A$ are pairwise coprime;
  3. 3)
Likewise, one defines an admissible set of initial variables $A\subseteq\tilde{\mathbf{x}}_{0}$ in $(\tilde{\mathcal{C}},\tilde{R})$. Note that for $A=\mathbf{x}_{0}$, the conditions listed above are precisely the conditions of the Starfish lemma (see Proposition 2.9).
Proposition 4.2.
Let $(\mathcal{C},R)$ and $(\tilde{\mathcal{C}},\tilde{R})$ be cluster structures of geometric type in normal Noetherian domains $R$ and $\tilde{R}$ with fixed initial extended clusters $\mathbf{x}_{0}$ and $\tilde{\mathbf{x}}_{0}$. Assume the following:
  1. 1)
  2. 2)
    The initial extended cluster $\tilde{\mathbf{x}}_{0}$ is admissible;
  3. 3)
    The set of marked variables $\mathbf{x}_{0}(\kappa)$ is admissible;
  4. 4)
    Each nonmarked cluster variable in $\mathbf{x}_{0}$ is regular and is coprime with each marked variable.
Then $\mathbf{x}_{0}$ is admissible, $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$.
Proposition 4.3.
Let $R$, $\tilde{R}$ and $\hat{R}$ be normal Noetherian domains, $(\mathcal{C},R)$, $(\tilde{\mathcal{C}},\tilde{R})$ and $(\hat{\mathcal{C}},\hat{R})$ be cluster structures of geometric type in $R$, $\tilde{R}$ and $\hat{R}$, with fixed initial extended clusters $\mathbf{x}_{0}$, $\tilde{\mathbf{x}}_{0}$ and $\hat{\mathbf{x}}_{0}$. Assume the following:
  1. 1)
    There are markers $\kappa$ and $\kappa^{\prime}$ that make $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ and $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$ related triples, as well as a pair of complementary birational quasi-isomorphisms $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$ and $\mathcal{Q}^{\prime}:(\mathcal{C},R)\rightarrow(\hat{\mathcal{C}},\hat{R})$;
  2. 2)
  3. 3)
    Every $x_{i}\in\mathbf{x}_{0}(\kappa)$ is coprime with every $x_{j}\in\mathbf{x}_{0}(\kappa^{\prime})$;
  4. 4)
    The initial extended cluster $\mathbf{x}_{0}$ is regular.
Then $\mathbf{x}_{0}$ is admissible, as well as $\bar{\mathcal{A}}(\hat{\mathcal{C}})\subseteq\hat{R}$, $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$.
The next two results are analogues of Proposition 4.2 and Proposition 4.3 in case the given rings are UFD’s. It is often desirable to show that the initial variables are irreducible; this allows, in turn, to drop Condition 3 in the previous proposition.
Definition 4.4.
  1. 1)
    All variables in $A$ are regular and irreducible;
  2. 2)
Proposition 4.5.
Let $(\mathcal{C},R)$ and $(\tilde{\mathcal{C}},\tilde{R})$ be cluster structures of geometric type in factorial Noetherian $\mathbb{C}$-algebras with no zero divisors and such that $R^{\times}=\tilde{R}^{\times}=\mathbb{C}^{\times}$. Let $\mathbf{x}_{0}$ and $\tilde{\mathbf{x}}_{0}$ be fixed initial extended clusters. Assume the following:
  1. 1)
  2. 2)
    The initial extended cluster $\tilde{\mathbf{x}}_{0}$ is i-admissible;
  3. 3)
    The set of marked variables $\mathbf{x}_{0}(\kappa)$ is i-admissible;
  4. 4)
    Each nonmarked cluster variable in $\mathbf{x}_{0}$ is regular and is coprime with each marked variable.
Then $\mathbf{x}_{0}$ is i-admissible, as well as $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})\subseteq R$.
Proposition 4.6.
Let $R$, $\tilde{R}$ and $\hat{R}$ be factorial Noetherian $\mathbb{C}$-algebras with no zero divisors and such that $R^{\times}=\tilde{R}^{\times}=\hat{R}^{\times}=\mathbb{C}^{\times}$, $(\mathcal{C},R)$, $(\tilde{\mathcal{C}},\tilde{R})$ and $(\hat{\mathcal{C}},\hat{R})$ be cluster structures of geometric type in $R$, $\tilde{R}$ and $\hat{R}$, with fixed initial extended clusters $\mathbf{x}_{0}$, $\tilde{\mathbf{x}}_{0}$ and $\hat{\mathbf{x}}_{0}$. Assume the following:
  1. 1)
    There are markers $\kappa$ and $\kappa^{\prime}$ that make $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ and $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$ related triples, as well as a pair of complementary birational quasi-isomorphisms $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$ and $\mathcal{Q}^{\prime}:(\mathcal{C},R)\rightarrow(\hat{\mathcal{C}},\hat{R})$;
  2. 2)
    The initial extended clusters $\tilde{\mathbf{x}}_{0}$ and $\hat{\mathbf{x}}_{0}$ are i-admissible;
  3. 3)
    The initial extended cluster $\mathbf{x}_{0}$ is regular.
Then $\mathbf{x}_{0}$ is i-admissible, as well as $\bar{\mathcal{A}}_{\mathbb{C}}(\hat{\mathcal{C}})\subseteq\hat{R}$, $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})\subseteq R$.
Remark 4.7.
The coordinate ring $\mathbb{C}[G]$ of a connected simply connected simple complex algebraic group $G$ is a factorial domain such that $(\mathbb{C}[G])^{\times}=\mathbb{C}^{\times}$ (see, for instance, [21]). Thus Propositions 4.5 and 4.6 cover the needs of the Gekhtman-Shapiro-Vainshtein program [11, 14] on the construction of multiple generalized cluster structures in $\mathbb{C}[G]$ for a given group $G$. At the cost of conciseness, both propositions can be extended in various ways. For instance, one can require that $R^{\times}$ and $\tilde{R}^{\times}$ are groups of Laurent monomials in the invertible nonmarked frozen variables with coefficients in $\mathbb{C}^{\times}$ (thus covering the case of $\mathbb{C}[\operatorname{GL}_{n}]$).
Remark 4.8.
In the definition of an i-admissible set of variables, we include the requirement that the frozen variables are also irreducible. This is only a matter of convenience: we tailor the statement towards our applications (see Remark 5.7). If the frozen variables in $\tilde{R}$ are not irreducible, they may not be irreducible in $R$, but the statements of the above results are still valid.
The next propositions concern the reverse inclusion $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})\supseteq R$ and are based on the result on upper bounds (see Propositions 2.7) in case there are birational quasi-isomorphisms. These results were present implicitly in the proof of [10, Theorem 3.11]. For the sake of completeness, we record them in the setup of the paper.
Denote by $\mathbf{x}_{i}^{\prime}$ the mutation of the initial extended cluster $\mathbf{x}_{0}$ in direction $i\in[1,N]$.
Proposition 4.9.
  1. 1)
  2. 2)
    The matrices $B_{0}$ and $\tilde{B}_{0}$ are of full rank;
  3. 3)
    The reverse inclusion $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\supseteq\tilde{R}$ holds;
  4. 4)
    For each $i\in\mathcal{I}(\kappa)$, $\mathcal{L}(\mathbf{x}_{i}^{\prime})\supseteq R$.
Then $\bar{\mathcal{A}}(\mathcal{C})\supseteq R$.
Remark 4.10.
Proposition 4.11.
Let $R$, $\tilde{R}$ and $\hat{R}$ be normal Noetherian domains, $(\mathcal{C},R)$, $(\tilde{\mathcal{C}},\tilde{R})$ and $(\hat{\mathcal{C}},\hat{R})$ be cluster structures of geometric type in $R$, $\tilde{R}$ and $\hat{R}$, with fixed initial extended seeds $(\mathbf{x}_{0},B_{0})$, $(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$ and $(\hat{\mathbf{x}}_{0},\hat{B}_{0})$. Assume the following:
  1. 1)
    There are markers $\kappa$ and $\kappa^{\prime}$ that make $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ and $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$ related triples, as well as a pair of complementary birational quasi-isomorphisms $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$ and $\mathcal{Q}^{\prime}:(\mathcal{C},R)\rightarrow(\hat{\mathcal{C}},\hat{R})$;
  2. 2)
    The matrices $\tilde{B}_{0}$, $\hat{B}_{0}$ and $B_{0}$ are of full rank;
  3. 3)
    The inclusions $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\supseteq\tilde{R}$ and $\bar{\mathcal{A}}(\hat{\mathcal{C}})\supseteq\hat{R}$ hold.
Then $\bar{\mathcal{A}}(\mathcal{C})\supseteq R$.
Remark 4.12.
None of the proofs of the above results uses a specific form of the mutation relation. The proofs rely on the following facts: 1) Starfish lemma (implicitly on the Laurent phenomenon); 2) If $\mathbf{x}_{0}$ is regular, then $x^{\prime}\in R[x^{\pm 1}]$ for every cluster variable $x\in\mathbf{x}_{0}$; 3) The equality $\bar{\mathcal{A}}(\mathcal{C})=\bar{\mathcal{A}}(\mathcal{N})$ for a star-shaped nerve $\mathcal{N}$ in $\mathcal{C}$. Thus the above results hold for more general cluster structures in which the mentioned facts are true. In particular, if Remark 3.10 is taken into account, the above results hold for generalized cluster structures as defined in [3, 15].
5 Single birational quasi-isomorphism
The setup for this section is as follows. Let $R$ and $\tilde{R}$ be normal Noetherian domains. Assume that $(\mathcal{C},R)$ and $(\tilde{\mathcal{C}},\tilde{R})$ are cluster structures of geometric type in $R$ and $\tilde{R}$ with fixed initial extended seeds $(\mathbf{x}_{0},B_{0})$ and $(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$, and such that $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ is a related triple. We also assume there is a birational quasi-isomorphism $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$. The objective of this section is to prove results that involve the use of a single birational quasi-isomorphism; that is, Proposition 4.2, Proposition 4.5 and Proposition 4.9, along with results of lesser importance. We note that, according to Definition 3.11, the marked variables $\mathbf{x}_{0}(\kappa)$ and $\tilde{\mathbf{x}}_{0}(\kappa)$ are assumed to be regular.